# Rethinking Graph Backdoor Attacks: A Distribution-Preserving Perspective

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have shown remarkable performance in various tasks. However, recent works reveal that GNNs are vulnerable to backdoor attacks. Generally, backdoor attack poisons the graph by attaching backdoor triggers and the target class label to a set of nodes in the training graph. A GNN trained on the poisoned graph will then be misled to predict test nodes attached with trigger to the target class. Despite their effectiveness, our empirical analysis shows that triggers generated by existing methods tend to be out-of-distribution (OOD), which significantly differ from the clean data. Hence, these injected triggers can be easily detected and pruned with widely used outlier detection methods in real-world applications. Therefore, in this paper, we study a novel problem of unnoticeable graph backdoor attacks with in-distribution (ID) triggers. To generate ID triggers, we introduce an OOD detector in conjunction with an adversarial learning strategy to generate the attributes of the triggers within distribution. To ensure a high attack success rate with ID triggers, we introduce novel modules designed to enhance trigger memorization by the victim model trained on poisoned graph. Extensive experiments on real-world datasets demonstrate the effectiveness of the proposed method in generating in distribution triggers that can by-pass various defense strategies while maintaining a high attack success rate.
Backdoor Attack, Graph Neural Networks
ccs: Computing methodologies Machine learningconference: the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph-structured data is pervasive in real world, such as social networks (Fan et al., 2019), molecular structures (Mansimov et al., 2019), and knowledge graphs (Liu et al., 2022b). With the growing interest in learning from graphs, Graph Neural Networks (GNNs), which have shown great ability in node representation learning on graphs, have become increasingly prominent. Generally, GNNs adopt the message-passing mechanism, which update a node’s representation by recursive propagation and aggregation of information from a node’s neighbors. The learned node representations preserve both node attributes and local graph structure information, which can benefit a range of downstream tasks, such as node classification (Kipf and Welling, 2017; Veličković et al., 2018; Hamilton et al., 2018), graph classification (Xu et al., 2019), and link prediction (Zhang and Chen, 2018).
Though GNNs have achieved remarkable performance across various applications, recent studies (Dai et al., 2023; Xi et al., 2021; Zhang et al., 2021a) have shown that they are vulnerable to backdoor attacks. Generally, backdoor attacks generate and attach backdoor triggers to a selected group of nodes, known as target nodes, and assign target nodes a target class. Triggers are typically a node or a subgraph and can either be predefined or generated by a trigger generator. When a GNN model is trained on a dataset poisoned with these triggers, it learns to associate the presence of the trigger with the target class. Consequently, during inference, the backdoored model will misclassify test nodes attached with the trigger to the target class, while maintain high prediction accuracy on clean nodes, i.e., nodes without triggers attached. Backdoor attacks on graphs pose a significant threat to the adoption of GNNs in real-world, especially on high-stake scenarios such as banking systems and cybersecurity. For example, an adversary could inject backdoor triggers to the training data for fraud detection in transaction networks, and bypass the detection of a model trained on such poisioned graph by disguising illegal behaviors with backdoor triggers.
Hence, graph backdoor attack is attracting increasing attention and several initial efforts have been taken (Zhang et al., 2021a; Xi et al., 2021; Dai et al., 2023). For example, SBA (Zhang et al., 2021a) conducts pioneering research on graph backdoor attacks. It adopts randomly generated graphs as triggers. Building on this work, GTA (Xi et al., 2021) adopts a backdoor trigger generator to generate more powerful sample-specific triggers to improve the attack success rate. Dai et al. (Dai et al., 2023) shows that the generated triggers in previous work can be easily broken by pruning edges linking nodes with low cosine feature similarity. To alleviate this issue, they propose UGBA, which adopts an unnoticeable constraint to make the triggers and the target nodes to have large cosine similarity of features.
@@
  • We design a novel graph backdoor attack framework, which can generate in-distribution triggers that can bypass outlier detection and achieve high attack success rate; and
@@ 2.1. Graph Neural Networks
With the increasing need for learning on graph structured data, Graph Neural Networks (GNNs), which have shown great power in modeling graphs, are developing rapidly in recent years (Dai et al., 2022; Zhao et al., 2021; Zhou et al., 2022). Most GNN variants operate under the message-passing framework, which integrates pattern extraction and interaction modeling across each layer (Hamilton et al., 2018; Zhang and Chen, 2018; Kipf and Welling, 2017). Essentially, GNNs handle messages derived from node representations, propagating these messages through various message-passing mechanisms to enhance node representations. These refined representations are subsequently applied to downstream tasks With the evolution of GNN technology, numerous advancements have been made to augment their performance and application scope. Innovations in self-supervised learning techniques for GNNs aim to lessen the dependency on annotated data (You et al., 2021; Lin et al., 2023; Zhu et al., 2020). Additionally, significant strides have been made in enhancing the fairness (Zhu et al., 2023; Dai and Wang, 2021), robustness (Dai et al., 2022) and interpretability of GNN frameworks (Zhang et al., 2021b). Furthermore, specialized GNN architectures have been developed to address the unique challenges presented by heterophilic graphs (Xu et al., 2022a), broadening the potential use cases of GNNs in complex networked systems.
@@ 2.2. Backdoor Attacks on Graph
Backdoor attacks have been widely studied in image domain (Li et al., 2022; Gu et al., 2019; Chen et al., 2017). Initial work directly poison training samples (Liu et al., 2020; Chen et al., 2017). Others have explored the invisibility of triggers (Li et al., 2021; Doan et al., 2021). Besides, the hidden backdoor could also be embedded through transfer learning (Kurita et al., 2020), modifying model parameters (Chen et al., 2021), and adding extra malicious modules (Tang et al., 2020). Recent studies have begun to delve into backdoor attacks on GNNs, focusing on a strategy distinct from the more prevalent poisoning and evasion attacks. Backdoor attacks involve injecting malicious triggers in the training data, which cause the model to make incorrect predictions when these triggers are presented in test samples. This form of attack subtly manipulates the training phase of a model, ensuring it performs as expected under regular conditions but fails in the presence of trigger-embedded inputs. Among the pioneering efforts, SBA (Zhang et al., 2021a) introduced a method for injecting universal triggers into training samples through a subgraph-based approach. However, the attack success rate is poor. GTA (Xi et al., 2021) furthered this by developing a technique for generating adaptive triggers, customizing perturbations for individual samples to enhance attack effectiveness. In UGBA (Dai et al., 2023), an algorithm for selecting poisoned nodes is introduced to optimize the utilization of the attack budget. Additionally, an adaptive trigger generator is employed to create triggers that demonstrate a high cosine similarity to the target node. While GTA and UGBA achieve a high attack success rate, the generated triggers tend to be outliers. This is because it is more efficient for a victim model to associate outlier triggers with the target class, leading the unconstrained trigger generator to exploit this shortcut for a higher attack success rate.
@@ 3.1. Notations
We denote an attributed graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\{v_{1},\dots,v_{N}\}$ represents the set of $N$ nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, and $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ denotes the set of node attributes, with $\mathbf{x}_{i}$ being the attribute of node $v_{i}$. The adjacency matrix of the graph $\mathcal{G}$ is denoted as $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected; otherwise, $A_{ij}=0$. In this paper, we concentrate on backdoor attack for semi-supervised node classification task within the inductive setting. Specifically, the training graph $\mathcal{G}$ includes a small subset of labeled nodes $\mathcal{V}_{L}\subseteq\mathcal{V}$ with labels as $\mathcal{Y}_{L}=\left\{y_{1},\dots,y_{N_{L}}\right\}$. The remaining nodes of $\mathcal{G}$ are unlabeled, denoted as $\mathcal{V}_{U}$. The test nodes, denoted as $\mathcal{V}_{T}$, are not part of the training graph $\mathcal{G}$, i.e., $\mathcal{V}_{T}\cap\mathcal{V}=\emptyset$. We aim to add backdoor triggers within budget to the training graph such that a GNN model trained on the backdoored graph will be fooled to give targeted label for test nodes attached with triggers.
@@ 3.2. Preliminaries of Graph Backdoor Attacks
Evaluation Setting: Given $\mathcal{V}_{P}\subseteq\mathcal{V}$ as a set of poisoned node, we attach the generated trigger $g_{i}=(\mathbf{X}^{g}_{i},\mathbf{A}^{g}_{i})$ to the node $v_{i}\in\mathcal{V}_{P}$ and assign $\mathcal{V}_{P}$ with target class $y_{t}$ to form the backdoored dataset. The victim model is then trained on this dataset. During inference, triggers generated by $f_{g}$ are attached to test nodes $v_{i}\in\mathcal{V}_{T}$ to evaluate the attack performance.
@@ 3.3. Outlier Issues of Graph Backdoor Attacks
To show that the triggers generated by existing graph backdoor attack methods are outliers, we conduct analysis on Pubmed dataset (Sen et al., 2008). We first adopt existing backdoor attack algorithms to add backdoors to the graph under the semi-supervised setting. We then apply Principal Component Analysis (PCA) to reduce the dimensionality of node attributes for both clean nodes and triggers’ nodes, and visualize them in a 2-dimensional space as shown in Figure 1, where the blue and the red dots denote the clean node and the generated triggers, respectively. From the figure, it is obvious that the generated triggers of GTA (Xi et al., 2021) and UGBA (Dai et al., 2023) are very different and far from the clean data distribution, showing that the triggers generated by many existing algorithms are outliers.
@@ 4.2.1. Enhancing Memorization of Triggers
(5) $\mathcal{L}_{E}=-\sum_{v_{i}\in\mathcal{V}}\sum_{v_{j}\in\mathcal{V}}S(z_{s}(\tilde{v}_{i}),z_{s}(\tilde{v}_{j}))+$
$\sum_{v_{i}\in\mathcal{V}}\sum_{v_{j}\in\mathcal{V}_{t}}-\log(\frac{S(z_{s}(\tilde{v}_{i}),z_{s}(v_{j}))}{S(z_{s}(\tilde{v}_{i}),z_{s}(v_{j}))+\sum_{v_{k}\in{\mathcal{V}_{L}}\backslash\mathcal{V}_{t}}S(z_{s}(\tilde{v}_{i}),z_{s}(v_{k}))}),$
@@ 4.2.2. Enhancing Attack Effectiveness against Unseen Targets
(6) $\mathcal{L}_{T}=\sum_{v_{i}\in\mathcal{V}}w_{i}\cdot l\left(f_{s}\left(\tilde{v}_{i}\right),y_{t}\right),$
@@ 5.1.2. Compared Methods
We compare DPGBA with representative and state-of-the-art graph backdoor attack methods, including UGBA (Dai et al., 2023), GTA (Xi et al., 2021), SBA-Samp (Zhang et al., 2021a) and its variant SBA-Gen. More details of these compared methods can be found in Appendix H. For a fair comparison, hyperparameters of all the attack methods are tuned based on the performance of the validation set.
@@
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA Ours
@@ Appendix B REPRESENTATIVE NODES SELECTION
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA Ours
@@
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA Ours
@@
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA Ours
@@ Appendix D ADDITIONAL RELATED WORKS
Graph outlier detection is a critical task in machine learning, involving the identification of anomalous nodes within a graph. Unlike traditional outlier detection on tabular or time-series data, graph outlier detection presents unique challenges due to the rich information inherent in graph structures and the computational complexity of training with complex machine learning models. The emergence of deep learning techniques has revolutionized outlier detection, shifting from traditional methods to neural network approaches (Ma et al., 2023). One popular neural network architecture for this task is the autoencoder (AE) (Kingma and Welling, 2022), which learns to reconstruct the original data and identifies outliers based on reconstruction errors. This unsupervised learning approach makes AEs effective for detecting outliers without the need for labeled data. Furthermore, graph neural networks (GNNs) have demonstrated superior performance in capturing complex patterns within graph data, considering both node attributes and graph structure. GNNs encode representations for each node, enabling effective outlier detection. Notably, GNNs can be combined with AEs (Ding et al., 2019; Fan et al., 2020; Kipf and Welling, 2016; Dai and Chen, 2022; Bandyopadhyay et al., 2018; Xu et al., 2022b), leveraging the strengths of both approaches for more robust outlier detection in graph data.

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have shown remarkable performance in various tasks. However, recent works reveal that GNNs are vulnerable to backdoor attacks. Generally, backdoor attack poisons the graph by attaching backdoor triggers and the target class label to a set of nodes in the training graph. A GNN trained on the poisoned graph will then be misled to predict test nodes attached with trigger to the target class. Despite their effectiveness, our empirical analysis shows that triggers generated by existing methods tend to be out-of-distribution (OOD), which significantly differ from the clean data. Hence, these injected triggers can be easily detected and pruned with widely used outlier detection methods in real-world applications. Therefore, in this paper, we study a novel problem of unnoticeable graph backdoor attacks with in-distribution (ID) triggers. To generate ID triggers, we introduce an OOD detector in conjunction with an adversarial learning strategy to generate the attributes of the triggers within distribution. To ensure a high attack success rate with ID triggers, we introduce novel modules designed to enhance trigger memorization by the victim model trained on poisoned graph. Extensive experiments on real-world datasets demonstrate the effectiveness of the proposed method in generating in distribution triggers that can bypass various defense strategies while maintaining a high attack success rate. Our code is available at: https://github.com/zzwjames/DPGBA.
Backdoor Attack; Graph Neural Networks
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671910isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph-structured data is pervasive in real world, such as social networks (Fan et al., 2019), molecular structures (Mansimov et al., 2019), and knowledge graphs (Liu et al., 2022b). With the growing interest in learning from graphs, Graph Neural Networks (GNNs), which have shown great ability in node representation learning on graphs, have become increasingly prominent. Generally, GNNs adopt the message-passing mechanism, which update a node’s representation by recursive propagation and aggregation of information from a node’s neighbors. The learned node representations preserve both node attributes and local graph structure information, which can benefit a range of downstream tasks, such as node classification (Kipf and Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017), graph classification (Xu et al., 2019), and link prediction (Zhang and Chen, 2018).
Though GNNs have achieved remarkable performance across various applications, recent studies (Dai et al., 2023; Xi et al., 2021; Zhang et al., 2021) have shown that they are vulnerable to backdoor attacks. Generally, backdoor attacks generate and attach backdoor triggers to a selected group of nodes, known as target nodes, and assign target nodes a target class. Triggers are typically a node or a subgraph and can either be predefined or generated by a trigger generator. When a GNN model is trained on a dataset poisoned with these triggers, it learns to associate the presence of the trigger with the target class. Consequently, during inference, the backdoored model will misclassify test nodes attached with the trigger to the target class, while maintain high prediction accuracy on clean nodes, i.e., nodes without triggers attached. Backdoor attacks on graphs pose a significant threat to the adoption of GNNs in real-world, especially on high-stake scenarios such as banking systems and cybersecurity. For example, an adversary could inject backdoor triggers to the training data for fraud detection in transaction networks, and bypass the detection of a model trained on such poisioned graph by disguising illegal behaviors with backdoor triggers.
Hence, graph backdoor attack is attracting increasing attention and several initial efforts have been taken (Zhang et al., 2021; Xi et al., 2021; Dai et al., 2023). For example, SBA (Zhang et al., 2021) conducts pioneering research on graph backdoor attacks. It adopts randomly generated graphs as triggers. Building on this work, GTA (Xi et al., 2021) adopts a backdoor trigger generator to generate more powerful sample-specific triggers to improve the attack success rate. Dai et al. (Dai et al., 2023) shows that the generated triggers in previous work can be easily broken by pruning edges linking nodes with low cosine feature similarity. To alleviate this issue, they propose UGBA, which adopts an unnoticeable constraint to make the triggers and the target nodes to have large cosine similarity of features.
@@
  • We design a novel graph backdoor attack framework, which can generate in-distribution triggers that can bypass outlier detection and achieve high attack success rate;
@@ 2.1. Graph Neural Networks
With the increasing need for learning on graph structured data, Graph Neural Networks (GNNs), which have shown great power in modeling graphs, are developing rapidly in recent years (Dai et al., 2022; Zhao et al., 2021; Zhou et al., 2022). Most GNN variants operate under the message-passing framework, which integrates pattern extraction and interaction modeling across each layer (Hamilton et al., 2017; Zhang and Chen, 2018; Kipf and Welling, 2017). Essentially, GNNs handle messages derived from node representations, propagating these messages through various message-passing mechanisms to enhance node representations. These refined representations are subsequently applied to downstream tasks With the evolution of GNN technology, numerous advancements have been made to augment their performance and application scope. Innovations in self-supervised learning techniques for GNNs aim to lessen the dependency on annotated data (You et al., 2020; Lin et al., 2023; Zhu et al., 2020; Ma et al., 2024; Xiao et al., 2024). Additionally, significant strides have been made in enhancing the fairness (Zhu et al., 2023; Dai and Wang, 2021), robustness (Dai et al., 2022) and interpretability of GNN frameworks (Zhang et al., 2022). Furthermore, specialized GNN architectures have been developed to address the unique challenges presented by heterophilic graphs (Xu et al., 2022a), broadening the potential use cases of GNNs in complex networked systems.
@@ 2.2. Backdoor Attacks on Graph
Backdoor attacks have been widely studied in image domain (Li et al., 2022; Gu et al., 2019; Chen et al., 2017). Initial work directly poison training samples (Liu et al., 2020; Chen et al., 2017). Others have explored the invisibility of triggers (Li et al., 2021; Doan et al., 2021). Besides, the hidden backdoor could also be embedded through transfer learning (Kurita et al., 2020), modifying model parameters (Chen et al., 2021), and adding extra malicious modules (Tang et al., 2020). Recent studies have begun to delve into backdoor attacks on GNNs, focusing on a strategy distinct from the more prevalent poisoning and evasion attacks. Backdoor attacks involve injecting malicious triggers in the training data, which cause the model to make incorrect predictions when these triggers are presented in test samples. This form of attack subtly manipulates the training phase of a model, ensuring it performs as expected under regular conditions but fails in the presence of trigger-embedded inputs. Among the pioneering efforts, SBA (Zhang et al., 2021) introduced a method for injecting universal triggers into training samples through a subgraph-based approach. However, the attack success rate is poor. GTA (Xi et al., 2021) furthered this by developing a technique for generating adaptive triggers, customizing perturbations for individual samples to enhance attack effectiveness. In UGBA (Dai et al., 2023), an algorithm for selecting poisoned nodes is introduced to optimize the utilization of the attack budget. Additionally, an adaptive trigger generator is employed to create triggers that demonstrate a high cosine similarity to the target node. While GTA and UGBA achieve a high attack success rate, the generated triggers tend to be outliers. This is because it is more efficient for a victim model to associate outlier triggers with the target class, leading the unconstrained trigger generator to exploit this shortcut for a higher attack success rate.
@@ 3.1. Notations
We denote an attributed graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\{v_{1},\dots,v_{N}\}$ represents the set of $N$ nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, and $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ denotes the set of node attributes, with $\mathbf{x}_{i}$ being the attribute of node $v_{i}$. The adjacency matrix of the graph $\mathcal{G}$ is denoted as $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected; otherwise, $A_{ij}=0$. In this paper, we concentrate on backdoor attack for semi-supervised node classification task within the inductive setting. Specifically, the training graph $\mathcal{G}$ includes a small subset of labeled nodes $\mathcal{V}_{L}\subseteq\mathcal{V}$ with labels as $\mathcal{Y}_{L}=\left\{y_{1},\dots,y_{N_{L}}\right\}$. The remaining nodes of $\mathcal{G}$ are unlabeled, denoted as $\mathcal{V}_{U}$. We denote $\mathcal{V}_{Tr}=\mathcal{V}_{L}\cup\mathcal{V}_{U}$ as the node set for the training graph. The test nodes, denoted as $\mathcal{V}_{T}$, are not part of the training graph $\mathcal{G}$, i.e., $\mathcal{V}_{T}\cap\mathcal{V}_{Tr}=\emptyset$. We aim to add backdoor triggers within budget to the training graph such that a GNN model trained on the backdoored graph will be fooled to give targeted label for test nodes attached with triggers.
@@ 3.2. Preliminaries of Graph Backdoor Attacks
Evaluation Setting: Given $\mathcal{V}_{P}\subseteq\mathcal{V}_{U}$ as a set of poisoned node, we attach the generated trigger $g_{i}=(\mathbf{X}^{g}_{i},\mathbf{A}^{g}_{i})$ to the node $v_{i}\in\mathcal{V}_{P}$ and assign $\mathcal{V}_{P}$ with target class $y_{t}$ to form the backdoored dataset. The victim model is then trained on this dataset. During inference, triggers generated by trigger generator $f_{g}$ are attached to test nodes $v_{i}\in\mathcal{V}_{T}$ to evaluate the attack performance.
@@ 3.3. Outlier Issues of Graph Backdoor Attacks
To show that the triggers generated by existing graph backdoor attack methods are outliers, we conduct analysis on Pubmed dataset (Sen et al., 2008). We first adopt existing backdoor attack algorithms to add backdoors to the graph under the semi-supervised setting. We then apply Principal Component Analysis (PCA) to reduce the dimensionality of node attributes for both clean nodes and triggers’ nodes, and visualize them in a 2-dimensional space as shown in Fig. 1, where the blue and the red dots denote the clean node and the generated triggers, respectively. From the figure, it is obvious that the generated triggers of GTA (Xi et al., 2021) and UGBA (Dai et al., 2023) are very different and far from the clean data distribution, showing that the triggers generated by many existing algorithms are outliers.
@@ 4.2.1. Enhancing Memorization of Triggers
(5) $\mathcal{L}_{E}=-\sum_{v_{i}\in\mathcal{V}_{U}}\sum_{v_{j}\in\mathcal{V}_{U}}S(z_{s}(\tilde{v}_{i}),z_{s}(\tilde{v}_{j}))+$
$\sum_{v_{i}\in\mathcal{V}_{U}}\sum_{v_{j}\in\mathcal{V}_{t}}-\log(\frac{S(z_{s}(\tilde{v}_{i}),z_{s}(v_{j}))}{S(z_{s}(\tilde{v}_{i}),z_{s}(v_{j}))+\sum_{v_{k}\in{\mathcal{V}_{L}}\backslash\mathcal{V}_{t}}S(z_{s}(\tilde{v}_{i}),z_{s}(v_{k}))}),$
@@ 4.2.2. Enhancing Attack Effectiveness against Unseen Targets
(6) $\mathcal{L}_{T}=\sum_{v_{i}\in\mathcal{V}_{U}}w_{i}\cdot l\left(f_{s}\left(\tilde{v}_{i}\right),y_{t}\right),$
@@ 5.1.2. Compared Methods
We compare DPGBA with representative and state-of-the-art graph backdoor attack methods, including UGBA (Dai et al., 2023), GTA (Xi et al., 2021), SBA-Samp (Zhang et al., 2021) and its variant SBA-Gen. More details of these compared methods can be found in Appendix H. For a fair comparison, hyperparameters of all the attack methods are tuned based on the performance of the validation set.
@@
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA DPGBA
@@ 6. Conclusion
ACKNOWLEDGMENTS
This material is based upon work supported by, or in part by, Army Research Office (ARO) under grant number W911NF-21-1- 0198, Department of Homeland Security (DHS) CINA under grant number E205949D, and Cisco Faculty Research Award.
@@ Appendix B REPRESENTATIVE NODES SELECTION
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA DPGBA
@@
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA DPGBA
@@
Datasets Defense Clean Graph SBA-Samp SBA-Gen GTA UGBA DPGBA
@@ Appendix D ADDITIONAL RELATED WORKS
Graph outlier detection is a critical task in machine learning, involving the identification of anomalous nodes within a graph. Unlike traditional outlier detection on tabular or time-series data, graph outlier detection presents unique challenges due to the rich information inherent in graph structures and the computational complexity of training with complex machine learning models. The emergence of deep learning techniques has revolutionized outlier detection, shifting from traditional methods to neural network approaches (Ma et al., 2021). One popular neural network architecture for this task is the autoencoder (AE) (Kingma and Welling, 2013), which learns to reconstruct the original data and identifies outliers based on reconstruction errors. This unsupervised learning approach makes AEs effective for detecting outliers without the need for labeled data. Furthermore, graph neural networks (GNNs) have demonstrated superior performance in capturing complex patterns within graph data, considering both node attributes and graph structure. GNNs encode representations for each node, enabling effective outlier detection. Notably, GNNs can be combined with AEs (Ding et al., 2019; Fan et al., 2020; Kipf and Welling, 2016; Dai and Chen, 2022; Bandyopadhyay et al., 2019; Xu et al., 2022b), leveraging the strengths of both approaches for more robust outlier detection in graph data.